### v1

@@ 5 Holographic Entanglement Entropy
A central consistency check of holographic duality is that the holographic entanglement entropy matches the area of the Ryu–Takayanagi (RT) surface in the bulk divided by $4G$ Ryu:2006bv ; Hubeny:2007xt . In this section, we first review the general expression for the entanglement entropy of the dual field theory in dS spacetime using the spherical partition function method Donnelly:2018bef ; Lewkowycz:2019xse ; Chen:2023eic ; Li:2020zjb , building on our previous analysis of the $T\bar{T}$ deformation of entanglement entropy Chang:2024voo  [^11] For further developments on entanglement entropy under the $T\bar{T}$ deformation, see Refs. Chen:2018eqk ; He:2019vzf ; Wang:2018jva ; He:2023xnb ; Apolo:2023ckr ; PhysRevD.108.126013 ; Jeong:2022jmp ; He:2023obo ; He:2023wko ; Deng:2023pjs ; Wang:2024jem ; Basu:2024enr ; Basu:2024xjq ; Basu:2025fsf ; Lai:2025thy . . We then compute the entanglement entropy and the corresponding RT surface areas for specific subregions, both within the static patch and in the extended regions. This allows us to test the consistency of the unified holographic framework proposed in this work.
@@ 5.1 Entanglement entropy in dS spacetime
This takes precisely the form of the pseudo entropy in the dS/CFT correspondence, reflecting the intrinsic non-unitarity of the dual field theory Doi:2022iyj ; Hikida:2022ltr . The real part of this expression has been confirmed holographically via the area of the RT surface in dS spacetime Narayan:2015vda ; Narayan:2022afv ; Narayan:2023zen .
@@ 5.3 Static spacetime: $r_{c}<\ell_{\rm dS}$
$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{\cos\tilde{\omega}_{c}\sin\phi_{a}}{\sqrt{1-\cos^{2}\tilde{\omega}_{c}\sin^{2}\phi_{a}}}\right).$ (84)
@@ 6 Conclusion and Discussion
In this paper, we employed the Wheeler–DeWitt equation and the Hamiltonian constraint to derive the dual field theory deformations corresponding to the inward motion of the boundary into the bulk in dS spacetime, considering spacelike and timelike boundaries separately. The inward motion of a spacelike boundary corresponds to the $T\bar{T}$ deformation, while that of the timelike boundary corresponds to $T\bar{T}+\Lambda_{2}$. By analyzing the energy spectrum, we found that for the $T\bar{T}$ deformation, the energy becomes complex when the deformation parameter reaches its critical value $\lambda_{0}$. By imposing the boundary condition $E_{\lambda_{0}}^{(T\bar{T})}=E_{\lambda_{0}}^{(T\bar{T}+\Lambda_{2})}$, the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations can be continuously connected, allowing the deformation parameter to continuously approach infinity. The transition from the $T\bar{T}$ to the $T\bar{T}+\Lambda_{2}$ deformation involves a change of the metric signature from Euclidean to Lorentzian. This coincides with with the transition of the boundary at $r=r_{c}$ from spacelike to timelike as it crosses the cosmological horizon. This observation motivates our proposal that the composite flow in dS holography is dual to the inward motion of the boundary at $r=r_{c}$ from the spacelike infinity in the extended spacetime toward the worldline of a static observer. Thus, our proposal provides a unified description of the two types of holographic models in dS holography.
The proposal is supported by the computation of holographic entanglement entropy. By employing the spherical partition function method, we first obtained the general entanglement entropy in the $T\bar{T}$-deformed dS/CFT. By separately considering the cases $r>\ell_{\mathrm{dS}}$ and $r<\ell_{\mathrm{dS}}$, we found two distinct results, corresponding to the pseudo entropy and the real entropy, respectively. These results precisely match the area of the RT surface on the bulk side.

### v2

@@ 5 Holographic Entanglement Entropy
A central consistency check of holographic duality is that the holographic entanglement entropy matches the area of the Ryu–Takayanagi (RT) surface in the bulk divided by $4G$ Ryu:2006bv ; Hubeny:2007xt . In this section, we first review the general expression for the entanglement entropy of the dual field theory in dS spacetime using the spherical partition function method Donnelly:2018bef ; Lewkowycz:2019xse ; Grieninger:2019zts ; Chen:2023eic ; Li:2020zjb , building on our previous analysis of the $T\bar{T}$ deformation of entanglement entropy Chang:2024voo  [^12] For further developments on entanglement entropy under the $T\bar{T}$ deformation, see Refs. Chen:2018eqk ; He:2019vzf ; Wang:2018jva ; He:2020qcs ; He:2022jyt ; He:2023xnb ; Apolo:2023ckr ; PhysRevD.108.126013 ; Jeong:2022jmp ; He:2023obo ; He:2023wko ; Deng:2023pjs ; Wang:2024jem ; Basu:2024enr ; Basu:2024xjq ; Basu:2025fsf ; Lai:2025thy ; Deng:2026xzv . . We then compute the entanglement entropy and the corresponding RT surface areas for specific subregions, both within the static patch and in the extended regions. This allows us to test the consistency of the unified holographic framework proposed in this work.
@@ 5.2 Extended spacetime: $r_{c}>\ell_{\rm dS}$
This expression takes precisely the form of the pseudo entropy in the dS/CFT correspondence, indicating that the density matrix of subsystem $A$ becomes non-Hermitian and thereby reflects the intrinsic non-unitarity of the dual field theory Doi:2022iyj ; Hikida:2022ltr . This expression has been confirmed holographically through complex extremal surfaces in dS spacetime Narayan:2015vda ; Narayan:2022afv ; Narayan:2023zen .
@@ 5.3 Static spacetime: $r_{c}<\ell_{\rm dS}$
$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{\cos\tilde{\omega}_{c}\sin\phi_{a}}{\sqrt{1-\cos^{2}\tilde{\omega}_{c}\sin^{2}\phi_{a}}}\right),$ (103)
where $\tilde{\omega}_{c}$ takes values in the range $(0,\frac{\pi}{2})$. In the case of dS static spacetime, the entanglement entropy of the subregion $A$ is real, which is consistent with the unitarity of the dual field theory. However, the entanglement entropy in this case takes the form of an arctangent function, which, unlike the usual logarithmic divergence of entanglement entropy in local field theories defined at an asymptotic boundary, remains finite. By analyzing the strong subadditivity and the boosted strong subadditivity of entanglement entropy, Refs. Miyaji:2015yva ; Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki ; Chang:2024voo showed that the dual field theory exhibits nonlocality in this setup.
For a unitary and local field theory, the entanglement entropy satisfies the strong subadditivity inequality
$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0\,.$ (104)
By considering infinitesimal boost transformations of subregions $A$ and $B$ within the dS spacetime context, one obtains the differential form of the boosted strong subadditivity Chang:2024voo , which can be written as
$S^{\prime\prime}(\Delta\phi)+\frac{1}{\sin\Delta\phi}S^{\prime}(\Delta\phi)\leq 0\,,\quad\Delta\phi=2\phi_{a}\,$, (105)
which is stronger than the standard strong subadditivity constraint.
Figure 7: The boosted strong subadditivity. Different colors represent different boundary positions, with the purple line indicating $\tilde{\omega}_{c}=0$, corresponding to the cosmological horizon $r=\ell_{\mathrm{dS}}$ in dS static spacetime, where the entanglement entropy of the dual field theory strongly violates the boosted strong subadditivity.
In the static spacetime case, the behavior of the boosted strong subadditivity for different boundary positions is illustrated in Fig. 7. It can be observed that in dS static patch holography, the entanglement entropy always violates the boosted strong subadditivity, indicating that the dual field theory defined on the timelike boundary is intrinsically nonlocal. Moreover, as the boundary approaches the cosmological horizon, the violation of the boosted strong subadditivity becomes increasingly pronounced. Finally, we further verify our entanglement entropy results by computing the corresponding geodesic lengths.
@@ 6 Conclusion and Discussion
In this paper, we employed the Wheeler–DeWitt equation and the Hamiltonian constraint to derive the dual field theory deformations corresponding to the inward motion of the boundary into the bulk in dS spacetime, considering spacelike and timelike boundaries separately. The inward motion of a spacelike boundary corresponds to the $T\bar{T}$ deformation, while that of the timelike boundary corresponds to $T\bar{T}+\Lambda_{2}$. By analyzing the energy spectrum, we found that for the $T\bar{T}$ deformation, the energy becomes complex when the deformation parameter reaches its critical value $\lambda_{0}$. By imposing the boundary condition $E_{\lambda_{0}}^{(T\bar{T})}=E_{\lambda_{0}}^{(T\bar{T}+\Lambda_{2})}$, the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations can be continuously connected, allowing the deformation parameter to continuously approach infinity. The transition from the $T\bar{T}$ to the $T\bar{T}+\Lambda_{2}$ deformation involves a change of the metric signature from Euclidean to Lorentzian. This coincides with with the transition of the boundary at $r=r_{c}$ from spacelike to timelike as it crosses the cosmological horizon. This observation motivates our proposal that the composite flow in dS holography is dual to the inward motion of the boundary at $r=r_{c}$ from the spacelike infinity in the extended spacetime toward the worldline of a static observer.
In the two classes of dS holographic models, the dual field theory is local but non-unitary when it is defined on a spacelike boundary, while it is unitary but nonlocal when it resides on a timelike boundary. The non-unitary and nonlocal properties can be indicated by the behavior of deformed entanglement entropy. In this paper, we proposed a possible approach toward a unified formulation of dS holography by exploiting the holographic duality of composite flows. The proposal is further supported by the computation of holographic entanglement entropy. By employing the spherical partition function method, we obtained a unified expression (85) for the entanglement entropy, which is valid in both static region $r\in(0,\ell_{\text{ds}})$ and extended spacetime $r\in(\ell_{\text{ds}},\infty)$. Specified to these two different regions $r>\ell_{\mathrm{dS}}$ and $r<\ell_{\mathrm{dS}}$, we found two distinct entanglment entropies, corresponding to the pseudo entropy (86) and the real entropy (87), respectively. These results precisely match the area of the RT surface on the bulk side.